OpenStudy (anonymous):

Please help. Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (anonymous):

make y the subject of the formulae for y + 2x = 9 then substitute it into the 1st equation, let me know if u have trouble.

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

y + 2x =9 y = 9 - 2x 3(9 - 2x) -2x = 11

OpenStudy (anonymous):

let me know if u understand and try and finish it :)

OpenStudy (anonymous):

3(9 - 2x) -2x = 11 27 - 6x - 2x = 11 -8x = - 16 x = 2

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

correct :) now solve for y

OpenStudy (anonymous):

y = 2(2) = 9 y = 9 - 4 y = 5

OpenStudy (anonymous):

correct :D congrates!

OpenStudy (anonymous):

Thank you so much! (:

OpenStudy (anonymous):

how will I check my answers?

OpenStudy (anonymous):

Input your value for y back in the original equation to check.